# Fractional Contribution of Dynamical and Geometric Phases in Quantum Evolution

### v1

@@ II The Main Result
where $H$ is the Hamiltonian of the system. Now, the action of the unitary operator on any state can be written as
$U(t)|\Psi(0)\rangle=\langle U(t)\rangle|\Psi(0)\rangle+\Delta U(t)|{\bar{\Psi}}(0)\rangle,$ (2)
where $|{\bar{\Psi}}(0)\rangle$ is orthogonal to $|\Psi(0)\rangle$, $\langle U(t)\rangle=\langle\Psi(0)|U(t)|\Psi(0)\rangle$ and $\Delta U(t)^{2}$ is the uncertainty in the unitary operator [20, 21, 22] defined as
$\Delta U(t)^{2}=$ (3)
$\langle\Psi(0)|U(t)^{\dagger}U(t)|\Psi(0)\rangle-\langle\Psi(0)|U(t)^{\dagger}|\Psi(0)\rangle\langle\Psi(0)|U(t)|\Psi(0)\rangle$
$=(1-|\langle\Psi(0)|\Psi(t)\rangle|^{2}).$
The Bargmann angle $S_{0}(t)$ or the shortest path connecting the initial state and the state at time $t$ is given by $|\langle\Psi(0)|\Psi(t)\rangle|=\cos\frac{S_{0}(t)}{2}$. Given its dynamic nature and fundamental role in quantifying quantum evolution, we call $S_{0}(t)$ as the time-dependent Bargmann angle. This also tracks the evolving geometric relationship between quantum states along a dynamical path, while distinguishing it from the static Bargmann angle between fixed states. Therefore, we have
$\langle U(t)\rangle=|\langle\Psi(0)|\Psi(t)\rangle|e^{i\Phi},$
where $\Phi=\arg\langle\Psi(0)|\Psi(t)\rangle$ is the Pancharatnam phase or the total phase. Therefore, the state at any later time can be expressed as
$|\Psi(t)\rangle=\cos\frac{S_{0}(t)}{2}e^{i\Phi}|\Psi(0)\rangle+\sin\frac{S_{0}(t)}{2}.|{\bar{\Psi}}(0)\rangle$ (4)
This shows that even though the quantum system lives in a $N$-dimensional Hilbert space, the time-evolution dynamics can be described within a two-dimensional subspace spanned by $\{|\Psi(0)\rangle,|{\bar{\Psi}}(0)\rangle\}$. This observation is of independent interest and can be exploited separately.
@@
where $|\chi(t)\rangle$ is the “reference-state” and defined as
@@
$|\chi(t)\rangle=\cos\frac{S_{0}(t)}{2}|\Psi(0)\rangle+\sin\frac{S_{0}(t)}{2}e^{-i\Phi}|{\bar{\Psi}}(0)\rangle.$ (8)
@@
Now, we show how the infinitesimal change in the dynamical phase and the geometric phase are connected to the infinitesimal change in the total phase during the quantum evolution. A quantitative measure of their relative contributions would provide a universal figure of merit to classify any quantum process. Using above definitions and Eq(5), our central result can be expressed as
$d\Phi_{D}=\cos^{2}\frac{S_{0}(t)}{2}d\Phi_{T}=|\langle\Psi(0)|\Psi(t)\rangle|^{2}d\Phi_{T},$ (9)
$d\Phi_{G}=\sin^{2}\frac{S_{0}(t)}{2}d\Phi_{T}=(1-|\langle\Psi(0)|\Psi(t)\rangle|^{2})d\Phi_{T}.$ (10)
@@
This shows that the fractional contributions of dynamical and geometric phases are completely governed by the instantaneous fidelity (Bargmann angle) and geometric separation between the initial quantum state and the state at time $t$. Our result also provides new insights to the instantaneous geometric “progress” of the quantum evolution relative to its starting point, making it particularly useful for studying quantum speed limits, optimal control, and the geometric structure of quantum dynamics. The relation $d\Phi_{G}=\sin^{2}\frac{S_{0}(t)}{2}d\Phi_{T}$ reveals a fundamental connection between the accumulation of geometric phase and the total phase during quantum evolution. This relation implies that the rate of geometric phase accumulation is modulated by the factor $\sin^{2}(S_{0}/2)$, which depends on the separation between the current state and the initial state. When $S_{0}$ is small (i.e., the state is close to the initial state), $\sin^{2}S_{0}/2\approx 0$. Thus, little geometric phase accumulates per unit total phase, indicating that most phase change is dynamical. When $S_{0}$ approaches $\pi$ (i.e., the state is nearly orthogonal to the initial state), $\sin^{2}S_{0}/2\approx 1$. Here, almost all total phase change converts into geometric phase, reflecting the dominant role of state-space curvature. This behavior underscores the idea that the geometry of quantum state space becomes more influential as the state evolves away from its starting point. The relation effectively measures how “efficiently” total phase changes are transformed into geometric phases, with the efficiency determined by the instantaneous state separation.
@@ III Geometric Phase Fraction
The geometric phase fraction is defined as the ratio of the geometric phase accumulation rate to the total phase accumulation rate during quantum evolution:
$f_{g}(t)=\frac{d\Phi_{G}}{d\Phi_{T}}=\sin^{2}\left(\frac{S_{0}(t)}{2}\right),$ (11)
where $S_{0}(t)$ is the instantaneous Bures angle between the initial state $|\Psi(0)\rangle$ and the state at time $t$. Its has following properties: (i) $f_{g}(t)\in[0,1]$, (ii) Monotonicity, i.e., generally increases with $S_{0}(t)$. The geometric phase fraction quantifies how “geometrically efficient” a quantum evolution is at a given moment. For example, if $f_{g}(t)\approx 0$, then the evolution is dominated by dynamical phase (state near initial position). If $f_{g}(t)\approx 1$, the evolution is dominated by geometric phase (state far from initial position). For other intermediate values, we have mixed contribution from both geometric and dynamical phases. Therefore, it provides a quantitative measure of how quantum evolution explores the curved geometry of state space, with implications for quantum information processing, fundamental physics, and quantum control theory. The concept highlights that not all quantum evolutions are equally “geometric,” and this geometric character has tangible consequences for quantum technologies and our understanding of quantum dynamics.
@@ IV Quantum Metric and Quantum Circuitousness
We know that when a quantum system evolves in time, it traces a path on the projective Hilbert space. The total distnace travelled by the quantum state can be measured using the Fubini-Study distance. The Fubini-Study metric provides a natural geometric measure of quantum evolution by quantifying the instantaneous separation between quantum states in the projective Hilbert space. For a quantum system evolving from time $|\Psi(t)\rangle$ to $|\Psi(t+dt)\rangle$, we can define the infinitesimal Fubini-Study distance [17, 18] as given by
@@ V Quantum Speed Limit and Geometric Phase
Here, we will prove a new quantum speed limit that depends on the geometric phase acquired by the system. From Eq(14), if we drop the second term, we can have $dS\geq dS_{0}$, and for a system evolving under a time-independent Hamiltonian, that will lead to the Mandelstam-Tamm [19] quantum speed limit bound, i.e.,
@@
where $S_{0}(T)=2\cos^{-1}|\langle\Psi(0)|\Psi(T)\rangle|$. Here, we will prove a new quantum speed limit that depends on the time-average of the geometric phase fraction and the geometric phase acquired during the time evolution. First, Note that we can express the infinitesimal Fubini-Study metric as
$dS^{2}(t)=dS_{0}t)^{2}+4d\Phi_{D}(t)d\Phi_{G}(t).$ (16)
Now, if we drop the first term and rearrange, we have
@@
This can be regarded as the geometric phase induced quantum speed limit, and it is independent of any QSL bounds known to date. The Mandelstamm-Tam bound on the QSL come from quantum dynamics, where as our bound comes from pure geometric phase. The geometric phase $\Phi_{G}(T)$ accumulates as the quantum state traverses an open path in the projective Hilbert space. Unlike the conventional QSL based on the Bures angle (which depends only on the initial and final states), the geometric phase is path-dependent. This means that if we have more winding in state space, that can tighten the QSL. By incorporating the geometric phase, the new QSL provides a precise figure of merit revealing the extent of non-geodesic evolution undertaken by the quantum state, quantifying how “circuitous” its path was.The bound suggests that the minimum time for a quantum process is determined by the ratio of the geometric path complexity to the speed of quantumm evolution (apart from other factors). The geometric complexity of paths is fundamentally different from conventional QSLs that depend only on the initial and final states.

### v2

@@ II The Main Result
where $H$ is the time-independent Hamiltonian of the system. Before we prove our main result, we need the following results.
Lemma 1: The action of an unitary operator $U$ on any state admits the orthogonal decomposition as given by
$U|\Psi\rangle=\langle U\rangle|\Psi\rangle+\Delta U|{\bar{\Psi}}(U)\rangle,$ (2)
where $|{\bar{\Psi}}(U)\rangle$ is orthogonal to $|\Psi\rangle$, $\langle U\rangle=\langle\Psi|U|\Psi\rangle$ and $\Delta U^{2}$ is the uncertainty in the unitary operator in the state $|\Psi\rangle$ [20, 21, 22] defined as
$\Delta U^{2}$ $=\langle U^{\dagger}U\rangle-\langle U^{\dagger}\rangle\langle U\rangle$ (3)
$=(1-|\langle\Psi|U|\Psi\rangle|^{2}).$
For proof, see Appendix.
Theorem (Unitary Evolution Decomposition): Let $U(t)$ be a unitary operator and $|\Psi(0)\rangle\in{\cal H}^{N}$ an initial quantum state. The time evolved state at any later time can be decomposed as
$|\Psi(t)\rangle=\cos\frac{S_{0}(t)}{2}e^{i\Phi}|\Psi(0)\rangle+\sin\frac{S_{0}(t)}{2}|{\bar{\Psi}}_{0}(t)\rangle,$ (4)
where $S_{0}(t)$ is the Bargmann angle $|\langle\Psi(0)|\Psi(t)\rangle|=\cos\frac{S_{0}(t)}{2}$ and $\Phi=\arg\langle\Psi(0)|\Psi(t)\rangle$ is the Pancharatnam phase or the total phase.
Proof: Let the unitary operator $U=U(t)$ and $|\Psi\rangle=|\Psi(0)\rangle$. Using the Lemma and denoting the orthogonal component $|{\bar{\Psi}}(U)\rangle=|{\bar{\Psi}}(U(t))\rangle=|{\bar{\Psi}}_{0}(t)\rangle$, we have $|\Psi(t)\rangle=\langle U(t)\rangle|\Psi(0)\rangle+\Delta U(t)|{\bar{\Psi}}_{0}(t)\rangle$. The average of the unitary $U(t)$ is $\langle\Psi(0)|U(t)|\Psi(0)\rangle=|\langle\Psi(0)|\Psi(t)\rangle|e^{i\Phi}=\cos\frac{S_{0}(t)}{2}e^{i\Phi}$ and the proof follows. Note that $|{\bar{\Psi}}(U)\rangle$ depends on the unitary via $U(t)|\Psi(0)\rangle$, which evolves in time (unless the evolution is trivial, e.g., $U(t)$ is a phase times identity).
The Bargmann angle $S_{0}(t)$ is also the shortest path connecting the initial state and the state at time $t$. Given its dynamic nature and fundamental role in quantifying quantum evolution, we call $S_{0}(t)$ as the time-dependent Bargmann angle. This also tracks the evolving geometric relationship between quantum states along a dynamical path, while distinguishing it from the static Bargmann angle between fixed states. This shows that even though the quantum system lives in a $N$-dimensional Hilbert space, the time-evolution dynamics can be described within a two-dimensional subspace spanned by $\{|\Psi(0)\rangle,|{\bar{\Psi}}_{0}(t)\rangle\}$. This observation is of independent interest and can be exploited separately.
@@
where $|\chi(t)\rangle$ is the “reference-state” and defined as [6, 7]
@@
$|\chi(t)\rangle=\cos\frac{S_{0}(t)}{2}|\Psi(0)\rangle+\sin\frac{S_{0}(t)}{2}e^{-i\Phi}|{\bar{\Psi}}_{0}(t)\rangle.$ (8)
@@
Now, we show how the infinitesimal change in the dynamical phase and the geometric phase are connected to the weighted sum and difference of infinitesimal change in the total phase and orthogonal phase, respectively, during the quantum evolution. A quantitative measure of their relative contributions would provide a universal figure of merit to classify any quantum process. Using above definitions and Eq. (5), our central result can be expressed as
$d\Phi_{D}=\cos^{2}\frac{S_{0}(t)}{2}d\Phi(t)+\sin^{2}\frac{S_{0}(t)}{2}d{\bar{\Phi}},$ (9)
$d\Phi_{G}=\sin^{2}\frac{S_{0}(t)}{2}d\Phi(t)-\sin^{2}\frac{S_{0}(t)}{2}d{\bar{\Phi}}.$
@@
Here, $d{\bar{\Phi}}$ is the infinitesimal phase coming from the evolution of the orthogonal component and ${\bar{\Phi}}=-i\int\langle{\bar{\Psi}_{0}(t)}|{\dot{\bar{\Psi}}_{0}(t)}\rangle~dt$. This shows that the fractional contributions of dynamical and geometric phases are completely governed by the instantaneous fidelity (Bargmann angle) and geometric separation between the initial quantum state and the state at time $t$. Our result also provides new insights to the instantaneous geometric “progress” of the quantum evolution relative to its starting point, making it particularly useful for studying quantum speed limits, optimal control, and the geometric structure of quantum dynamics.
Interpretation of the Dynamical Phase Relation: This relation shows that the total dynamical phase is a weighted average of two contributions: (i) $\cos^{2}(S_{0}/2)d\Phi$: Contribution from the total phase weighted by overlap with initial state $\cos^{2}(S_{0}/2)=|\langle\Psi(0)|\Psi(t)\rangle|^{2}$ which is the Bargmann angle. When state is near initial state ($S_{0}\approx 0$), this dominates. (ii) $\sin^{2}(S_{0}/2)d{\bar{\Phi}}$: Contribution from orthogonal component’s phase evolution weighted by $\sin^{2}(S_{0}/2)=1-|\langle\Psi(0)|\Psi(t)\rangle|^{2}$ which is infidelity. When state is far from initial state ($S_{0}\approx\pi$), this dominates.
Physical Interpretation of the Geometric Phase Relation: The relation $d\Phi_{G}=\sin^{2}(S_{0}(t)/2)(d\Phi-d{\bar{\Phi}})$ reveals that the infinitesimal geometric phase arises from a fundamental phase mismatch between the global evolution of the quantum state and the internal evolution of its orthogonal component, scaled by how far the system has explored Hilbert space from its initial configuration. Here, $d\Phi$ represents the infinitesimal Pancharatnam phase of the state, while $d{\bar{\Phi}}$ captures the infinitesimal phase accumulated by the orthogonal part $|{\bar{\Psi}}_{0}(t)\rangle$. The prefactor $\sin^{2}(S_{0}(t)/2)$, which equals the infidelity $1-|\langle\Psi(0)|\Psi(t)\rangle|^{2}$, acts as a geometric weight ensuring that the geometric phase only becomes significant when the state has substantially evolved away from its origin. This elegantly expresses that true geometric phase is not merely a global phase change, but a gauge-invariant measure of the anholonomy due to the curvature of the quantum state space, emerging from the differential rotation between the state’s components in a moving frame. When $S_{0}$ is small (i.e., the state is close to the initial state), $\sin^{2}S_{0}/2\approx 0$. Thus, little geometric phase accumulates. This behavior underscores the idea that the geometry of quantum state space becomes more influential as the state evolves away from its starting point. The relation effectively measures how “efficiently” the phase mismatch between the total phase changes and orthogonal phase is transformed into geometric phases, with the efficiency determined by the instantaneous state separation.
@@ III Geometric Phase Fraction
The geometric phase fraction is defined as the ratio of the geometric phase accumulation rate to the phase miss match rate during quantum evolution:
$f_{g}(t)=\frac{d\Phi_{G}}{(d\Phi-d{\bar{\Phi}})}=\sin^{2}\left(\frac{S_{0}(t)}{2}\right),$ (10)
where $S_{0}(t)$ is the instantaneous Bures angle between the initial state $|\Psi(0)\rangle$ and the state at time $t$. It has following properties: (i) $f_{g}(t)\in[0,1]$, (ii) Monotonicity, i.e., generally increases with $S_{0}(t)$ and (iii) This is also gauge invariant (both numerator and denominator shift equally under gauge transformations). The geometric phase fraction quantifies how “geometrically efficient” a quantum evolution is at a given moment. For example, if $f_{g}(t)\approx 0$, then the evolution is dominated by dynamical phase (state near initial position). If $f_{g}(t)\approx 1$, the evolution is dominated by geometric phase (state far from initial position). For other intermediate values, we have mixed contribution from both geometric and dynamical phases. Therefore, it provides a quantitative measure of how quantum evolution explores the curved geometry of state space, with implications for quantum information processing, fundamental physics, and quantum control theory. The concept highlights that not all quantum evolutions are equally “geometric,” and this geometric character has tangible consequences for quantum technologies and our understanding of quantum dynamics.
@@ IV Quantum Metric and Quantum Circuitousness
We know that when a quantum system evolves in time, it traces a path on the projective Hilbert space. The total distance travelled by the quantum state can be measured using the Fubini-Study distance. The Fubini-Study metric provides a natural geometric measure of quantum evolution by quantifying the instantaneous separation between quantum states in the projective Hilbert space. For a quantum system evolving from time $|\Psi(t)\rangle$ to $|\Psi(t+dt)\rangle$, we can define the infinitesimal Fubini-Study distance [17, 18] as given by
@@
To make connection to the Fubini-Study metric and quantum speed limit, we go to a controlled-rotating frame where we can transform the time-dependent orthogonal component to time-independent. In the orthogonal subspace, let $V(t)$ is a unitary that connects $|{\bar{\Psi}}_{0}(0)\rangle$ to $|{\bar{\Psi}}_{0}(t)\rangle$, i.e., $|{\bar{\Psi}}_{0}(t)\rangle=V(t)|{\bar{\Psi}}_{0}(0)\rangle$. In this frame, we can write the state at time $t$ as
$|\Psi(t)\rangle=\cos\frac{S_{0}(t)}{2}e^{i\Phi}|\Psi(0)\rangle+\sin\frac{S_{0}(t)}{2}|{\bar{\Psi}}_{0}(0)\rangle.$ (13)
@@ V Quantum Speed Limit and Geometric Phase
Here, we will prove a new quantum speed limit that depends on the geometric phase acquired by the system. From Eq. (14), if we drop the second term, we can have $dS\geq dS_{0}$, and for a system evolving under a time-independent Hamiltonian, that will lead to the Mandelstam-Tamm [19] quantum speed limit bound, i.e.,
@@
where $S_{0}(T)=2\cos^{-1}|\langle\Psi(0)|\Psi(T)\rangle|$. Here, we will prove a new quantum speed limit that depends on the time-average of the geometric phase fraction and the geometric phase acquired during the time evolution. First, note that in the controlled-rotating frame, we can express the infinitesimal Fubini-Study metric as
$dS^{2}(t)=dS_{0}(t)^{2}+4d\Phi_{D}(t)d\Phi_{G}(t).$ (16)
It may be noted that the contributions to the dynamical phase and the geometric phase can depend on the rotating frame [23]. However, we are not exploring this in detail here. We focus on the application of the metric structure in proving quantum speed limit. Now, if we drop the first term and rearrange, we have
@@
This can be regarded as the geometric phase induced quantum speed limit, and it is independent of any QSL bounds known to date. The Mandelstamm-Tam bound on the QSL come from quantum dynamics, where as our bound comes from pure geometric phase. The geometric phase $\Phi_{G}(T)$ accumulates as the quantum state traverses an open path in the projective Hilbert space. Unlike the conventional QSL based on the Bures angle (which depends only on the initial and final states), the geometric phase is path-dependent. This means that if we have more winding in state space, that can tighten the QSL. By incorporating the geometric phase, the new QSL provides a precise figure of merit revealing the extent of non-geodesic evolution undertaken by the quantum state, quantifying how “circuitous” its path was. The bound suggests that the minimum time for a quantum process is determined by the ratio of the geometric path complexity to the speed of quantum evolution (apart from other factors). The geometric complexity of paths is fundamentally different from conventional QSLs that depend only on the initial and final states.